# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v2

@@ 4 Practical Implications
where $\bar{\sigma}(.)$ is a non-linear function. We name the new attention mechanism (20) as active-attention. Our experiments with the active-attention in Figure 2 and Table 2 for both classification and time series forecasting tasks with a wide range of non-linear function $\bar{\sigma}(\cdot)$ demonstrate the favorable performance of active-attention over the standard attention mechanism.
@@
5 Experiments
In this section, we conduct numerical experiments to verify the theoretical results presented in Section 3, the favorable performance of the proposed active-attention mechanism (20) over standard attention mechanism, and the empirical benefits of quadratic gating over standard linear gating in language modeling.
Verification of theoretical results. We generate synthetic data based on the model described in equation (10). Details regarding the values of the true parameters and the training procedure can be found in the Appendix F.
@@ 5 Experiments
Performance of active-attention mechanism. We empirically demonstrate the effects of employing nonlinear activation functions in the proposed active-attention mechanism (20). Our experiments span large-scale image classification tasks on CIFAR-10 [18] and ImageNet [36], language modeling on WikiText-103 [23], and multivariate time series forecasting across 8 different benchmarks. We evaluate the impact of five commonly used activation functions, including ReLU [1], GELU [13], SiLU [9], Sigmoid, and Tanh [29]. Figures 2 (a) and (b) present the results of image classification using different activation functions in self-attention, with ViT [7] and CaiT [39] as the base models. For CIFAR-10, we employed the ViT-Tiny and CaiT-Tiny models, while for ImageNet, we utilized the ViT-Base and CaiT-Medium models. Figure 2 (c) displays the results of the large-scale language modeling task on WikiText-103, using the standard multi-head self-attention transformer [41]. Additionally, we tested various activation functions on the Performer model [5] as another backbone. Our findings show that the GELU and ReLU activation functions greatly improve performance compared to linear activation functions. This aligns with prior research, suggesting that these two activation functions are preferred in large-scale deep networks due to their ability to support more efficient and stable training.
Table 2 further evaluates the impact of different activation functions on transformer-based time-series forecasting models across eight forecasting tasks. In this experiment, we employ the state-of-the-art PatchTST model [28] and the standard self-attention transformer as the backbone. Unlike the results observed in Figure 2, in addition to GELU, we find that Tanh and Sigmoid functions also show prominent advantages over linear activation function. We hypothesize that this is due to the smoothing gradients provided by Tanh and Sigmoid, which may help in capturing subtle patterns in time-series data. Additionally, since these tasks tend to be smaller and more prone to overfitting, the saturation effects of Tanh and Sigmoid could serve as a regularization mechanism by limiting output ranges and avoiding extreme activations.
Quadratic gating versus linear gating. We performed experiments using GPT2 (124M) [33] MoE models on a dataset consisting of 10 billion tokens from FineWeb-Edu [30]. We focus on a GPT2 MoE model featuring 8 experts and a Top2 router with a quadratic gating network. In addition, we considered a linear gating MoE model and a dense GPT2 model for baseline comparisons. The hidden size of the experts is chosen so that all models have approximately the same number of activated parameters. Please refer to Appendix F for more details.
The tested models, along with the baselines, were evaluated on the HellaSwag benchmark. The GPT2-MoE model with a Top2 quadratic router achieved better performance than the dense model and the MoE model with linear gating, based on validation loss and HellaSwag accuracy.
  Model Val. Loss HellaSwag (%)
Dense 3.0211 30.74%
MoE (linear) 2.9928 29.95%
MoE (quadratic) 2.9712 32.11%
@@ 6 Discussion
In this paper, we first establish a link between the MoE framework and attention mechanisms. We introduce a formal attention-based approach for the MoE framework and show that quadratic gating in the MoE framework can be interpreted as a self-attention mechanism. Next, we carry out the convergence analysis of parameter and expert estimation under the MoE models with the quadratic polynomial gate and the quadratic monomial gate. Our theories indicate that experts formulated as neural networks with popular activation functions such as $\mathrm{ReLU}$ and $\tanh$ have faster estimation rates than linear experts. The insights from the theories lead to the new attention mechanism, named active-attention, where we replace the linear value matrix in the attention by non-linear value matrix. Through extensive empirical evaluation, we show the favorable performance of the proposed active-attention over standard attention in various tasks.
@@
Supplement to “Quadratic Gating Functions in Mixture of Experts: A Statistical Insight”
In this supplementary material, we first discuss related works to the quadratic gating MoE model in Appendix A. Then, we establish the expert estimation rates under the quadratic monomial gating MoE model in Appendix B. Subsequently, in Appendix C, we provide a convergence analysis for parameter and expert estimation under the quadratic gating mixture of linear experts. Full proofs for the theoretical results of Section 3 and Appendix C are presented in Appendix D. Next, we study the identifiability of the quadratic gating MoE in Appendix E. Lastly, we specify the experimental details for Figure 1 in Appendix F.
Appendix A Related Works
A more generalized version of quadratic MoE was first introduced as an alternative model for MoE, utilizing a distinct parametric structure in the gating network [42]. The proposed modified gating network is given by
$g_{i}({\bm{x}};\bm{\Theta}_{g})$ $=\frac{\pi_{i}~{}p({\bm{x}}\mid{\bm{\theta}}_{i})}{\sum_{j=1}^{N}\pi_{j}~{}p({\bm{x}}\mid{\bm{\theta}}_{j})},\quad i=1,\ldots,N,$ (21)
where $(\pi_{1},\ldots,\pi_{N})^{\top}\in\Delta^{N-1}$, and $\bm{\Theta}_{g}=\{(\pi_{i},{\bm{\theta}}_{i}),i=1,\ldots,N\}$ represents the learnable parameters, with each $p({\bm{x}}\mid{\bm{\theta}}_{i})$ being a density function from the exponential family. The gating function in equation (21) is a nonlinear variant of the softmax linear gating function. In particular, if we assume that $p({\bm{x}}\mid{\bm{\theta}}_{i})$ is a Gaussian density function with mean $\bm{\mu}_{i}\in\mathbb{R}^{d}$ and covariance matrix $\bm{\Sigma}_{i}\in\mathbb{R}^{d\times d}$, this leads to a specific form of quadratic softmax gating function. Here, it is assumed that the covariance matrices are positive definite. Observe that setting $\bm{\Sigma}_{i}=\bm{\Sigma}$ for all $i\in[N]$ reinstates the linear gating model.
The gating function $g_{i}({\bm{x}};\bm{\Theta}_{g})$ essentially models the posterior probability $\mathbb{P}(\zeta=i\mid{\bm{x}})$, indicating the likelihood that ${\bm{x}}$ is assigned to the partition associated with the $i$-th expert. Here, $\zeta\in\{1,\ldots,N\}$ is a latent gating variable that selects a particular expert. More precisely, the gating function defined in Equation (21) interprets this posterior probability when $\zeta$ follows a categorical distribution with parameters $(\pi_{1},\ldots,\pi_{N})^{\top}\in\Delta^{N-1}$, and conditioned on the event that $\zeta$ selects the $i$th expert, the distribution of ${\bm{x}}$ is modeled by a specific parametric distribution $p(~{}\cdot\mid{\bm{\theta}}_{i})$. Motivated by this interpretation of gating function, [21] proposed the Quadratically Gated Mixture of Experts, in which the parametric distribution $p(~{}\cdot\mid{\bm{\theta}}_{i})$ is assumed to follow a Gaussian density.
Appendix B Quadratic Monomial Gate
In this section, we proceed to streamline the analysis of the quadratic monomial gating based on the regression framework in equation (10). Due to the change of the gating function, the corresponding regression function is reformulated as follows:
$\tilde{f}_{G_{*}}({\bm{x}}):=\sum_{i=1}^{N^{*}}\frac{\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+c^{*}_{i})}{\sum_{j=1}^{N^{*}}\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{j}{\bm{x}}+c^{*}_{j})}\cdot h({\bm{x}},{\bm{\eta}}^{*}_{i}).$ (22)
In comparison with the quadratic polynomial gating, the first-degree monomial term $b^{\top}x$ has been removed from the scoring function. As a consequence, the least squares estimator under this setting also changes accordingly to
$\widetilde{G}_{n}:=\operatorname*{arg\,min}_{G\in\mathcal{G}_{N}(\Theta)}\sum_{i=1}^{n}\Big{(}y_{i}-\tilde{f}_{G}({\bm{x}}_{i})\Big{)}^{2}.$ (23)
Given the above estimator, we provide in Theorem B.1 the convergence rate of regression estimation $\tilde{f}_{\widetilde{G}_{n}}(\cdot)$ to the regression function $\tilde{f}_{G_{*}}(\cdot)$.
Theorem B.1 (Regression Estimation Rate).
$\|\tilde{f}_{\widetilde{G}_{n}}-\tilde{f}_{G_{*}}\|_{L_{2}(\mu)}=\mathcal{\widetilde{O}}_{P}(n^{-1/2}).$ (24)
See Appendix D.4 for the proof of Theorem B.1. It follows from the bound (24) that the regression estimation rate still remains parametric on the sample size, which matches that in Theorem 3.1 where we use the quadratic polynomial gating function in the MoE-type regression function.
Analogous to Section 3.1, we also derive a mono-strong identifiability condition in Definition B.2 to determine which expert functions will have faster estimation rates than others.
Definition B.2 (Mono-strong identifiability).
We say that an expert function $x\mapsto h({\bm{x}},{\bm{\eta}})$ is strongly identifiable if it is twice differentiable w.r.t its parameter ${\bm{\eta}}$, and if for any $k\geq 1$ and pair-wise different ${\bm{\eta}}_{1},\ldots,{\bm{\eta}}_{k}$, the following set
[…]
is linearly independent for almost every ${\bm{x}}$.
Example. It can be verified that the mono-strong identifiability condition holds for experts formulated as feed-forward neural networks with activation functions such as $\mathrm{ReLU}(\cdot)$ and $\tanh(\cdot)$. However, a linear expert fails to satisfy this condition.
$\mathcal{L}_{3}(G,G_{*}):=\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|^{2}+\|\Delta{\bm{\eta}}_{ij}\|^{2}\Big{]}$ $+\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|+\|\Delta{\bm{\eta}}_{ij}\|\Big{]}$
$+\sum_{j=1}^{N^{*}}\Big{|}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})-\exp(c_{j}^{*})\Big{|}.$ (25)
Now, we are ready to capture the parameter and expert estimation rates in Theorem B.3.
Theorem B.3.
Assume that the expert function $h({\bm{x}},{\bm{\eta}})$ is strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{G}_{N}(\Theta)$:
which together with Theorem 3.1 indicates that $\mathcal{L}_{3}(\widetilde{G}_{n},G_{*})=\mathcal{\widetilde{O}}_{P}(n^{-1/2})$.
Proof of Theorem B.3 is in Appendix D.5. A few comments regarding this theorem are in order: (i) The rates for estimating gating parameters ${\bm{A}}^{*}_{j}$ fitted by more than one atom, i.e. $|\mathcal{V}_{j}(\widetilde{G}_{n})|>1$, are significantly improved to be of order $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$. Those rates are much faster than their counterparts when using the quadratic polynomial gate, which stand at order $\widetilde{\mathcal{O}}_{P}(n^{-1/\bar{r}(|\mathcal{V}_{j}|)})$ (cf. Theorem 3.3). This rate acceleration is due to the disappearance of the interaction among gating parameters in equation (14) when using the quadratic monomial gate. Meanwhile, the estimation rates for expert parameters ${\bm{\eta}}^{*}_{j}$ remained unchanged at order $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$; (ii) Model parameters ${\bm{A}}^{*}_{j},{\bm{\eta}}^{*}_{j}$ fitted by exactly one atom, i.e. $|\mathcal{V}_{j}(\widetilde{G}_{n})|=1$, enjoy the parametric estimation rates of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$, which are comparable to their counterparts in Theorem 3.3.
Mono-weak identifiability of linear experts. Similar to the polynomial quadratic gating, the linear expert $h({\bm{x}},({\bm{\beta}}_{1},\beta_{0}))=({\bm{\beta}}_{1i})^{\top}{\bm{x}}+\beta_{0i}$ in the monomial quadratic gating setting also does not satisfy the mono-strong idenfiability condition. That violation leads to $\mathcal{O}(1/\log(n))$ rates of the parameters and experts under the quadratic monomial gating MoE. The proof for this result is similar to that of Theorem C.1 in Appendix C; therefore, it is omitted.
Appendix C Convergence Analysis for the Quadratic Gating Mixture of Linear Experts
In this appendix, we provide the convergence rates for parameter and expert estimation under the MoE model with the quadratic polynomial gate. Meanwhile, the analysis for the quadratic monomial gate can be done in a similar fashion.
As being mentioned in the main text, for the linear expert function $h({\bm{x}},({\bm{\beta}}_{1},\beta_{0}))=({\bm{\beta}}_{1i})^{\top}{\bm{x}}+\beta_{0i})$, where $({\bm{\beta}}_{1},\beta_{0})\in\mathbb{R}^{d}\times\mathbb{R}$, we observe that it violates the strong identifiability condition due to an interaction among parameters via the following partial differential equation:
$\frac{\partial^{2}F}{\partial b\partial\beta_{0}}({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\beta}}_{1},\beta_{0})=\frac{\partial F}{\partial{\bm{\beta}}_{1}}({\bm{x}};{\bm{A}}_{i},{\bm{b}}_{i},{\bm{\beta}}_{1},\beta_{0}),$ (26)
where we denote $F({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\beta}}_{1},\beta_{0}):=\exp({\bm{x}}^{\top}{\bm{A}}{\bm{x}}+{\bm{b}}^{\top}{\bm{x}})({\bm{\beta}}_{1}^{\top}{\bm{x}}+\beta_{0})$.
To capture the effects of such parameter interaction on the convergence of parameter estimation, let us design another Voronoi loss tailored to this setting. More specifically, we define for any $r\geq 1$ that
Theorem C.1.
Assume that the experts take the form ${\bm{\beta}}_{1}^{\top}{\bm{x}}+\beta_{0}$, then we achieve the following minimax lower bound of estimating $G_{*}$:
$\inf_{\overline{G}_{n}\in\mathcal{G}_{N}(\Theta)}\sup_{G\in\mathcal{G}_{N}(\Theta)\setminus\mathcal{G}_{N^{*}-1}(\Theta)}\mathbb{E}_{f_{G}}[\mathcal{L}_{2,r}(\overline{G}_{n},G)]\gtrsim n^{-1/2},$
Proof of Theorem C.1 is in Appendix D.3. A few remarks on the result of this theorem are in order. First, Theorem C.1 reveals that using linear experts make the estimation rates for all the parameters ${\bm{A}}^{*}_{i}$, ${\bm{b}}^{*}_{i}$, ${\bm{\beta}}^{*}_{1i}$ and $\beta^{*}_{0i}$ are slower than $\widetilde{\mathcal{O}}_{P}(n^{-1/2r})$ for any $r\geq 1$, and could be as slow as $\mathcal{O}_{P}(1/\log(n))$ owing to the interaction in equation (19). Second, we have that
Since the input space $\mathcal{X}$ is bounded, the rates for estimating linear experts $({\bm{\beta}}^{*}_{1j})^{\top}{\bm{x}}+\beta^{*}_{0j}$ could also be of order $\mathcal{O}_{P}(1/\log(n))$. Hence, combining with the result in Theorem 3.3, we deduce that the performance of a mixture of linear experts cannot compare to that of a mixture of non-linear experts in terms of the expert estimation problem. This observation totally aligns with the findings in [4].
Appendix D Proof of Theoretical Results

### v3

@@ 3.3 Practical Implications
where $\varphi$ is a non-linear activation function. We name the new attention mechanism (25) as active-attention. Our experiments with the active-attention in Figure 2 and Table 4 for both classification and time series forecasting tasks with a wide range of non-linear functions demonstrate the favorable performance of active-attention over the standard attention mechanism.
2. Benefits of quadratic monomial gating (8) over quadratic polynomial gating (7). The remarks after Theorems 3.3 and 3.7 indicate that the estimation rates of the gating parameters are independent of the amount of over-specification of the number of experts and much better than those of the polynomial gating parameters, which become very slow even when we only overspecify the model by a few experts. That theoretical advantage of the monomial gating over the polynomial gating confirms the benefits of the widely used linear embeddings of the keys and queries without bias terms in the attention in practice.
4 Related Work
Quadratic Gating MoE. Xu et al. (xu1994alternative, ) introduced an alternative model for MoE formulation, which incorporates a gating network with a distinct parametric structure. Their gating function is defined by
$\mathcal{G}({\bm{x}})_{i}$ $=\frac{\pi_{i}p({\bm{x}}\mid{\bm{\theta}}_{i})}{\sum_{j=1}^{N}\pi_{j}p({\bm{x}}\mid{\bm{\theta}}_{j})},\quad i=1,\ldots,N,$ (26)
where $(\pi_{1},\ldots,\pi_{N})^{\top}$ belongs to the simplex $\Delta^{N-1}$, and ${(\pi_{i},{\bm{\theta}}_{i}),i=1,\ldots,N}$ are learnable parameters. Here, each density $p({\bm{x}}\mid{\bm{\theta}}_{i})$ is from the exponential family. When specifically choosing Gaussian densities with mean $\bm{\mu}_{i}\in\mathbb{R}^{d}$ and covariance matrix $\bm{\Sigma}_{i}\in\mathbb{R}^{d\times d}$, the gating function naturally takes a quadratic form:
$\mathcal{G}({\bm{x}})_{i}=\frac{\exp\left(-\frac{1}{2}({\bm{x}}-\bm{\mu}_{i})^{\top}\bm{\Sigma}_{i}^{-1}({\bm{x}}-\bm{\mu}_{i})+\log\pi_{i}\right)}{\sum_{j=1}^{N}\exp\left(-\frac{1}{2}({\bm{x}}-\bm{\mu}_{j})^{\top}\bm{\Sigma}_{j}^{-1}({\bm{x}}-\bm{\mu}_{j})+\log\pi_{j}\right)}.$
This quadratic gating approach enables explicit analytic parameter updates within the Expectation Maximization (EM) algorithm, eliminating the need for iterative nonlinear optimization during training. As a result, the authors introduced a simplified, single-loop EM algorithm, significantly reducing computational complexity and improving convergence speed compared to traditional double-loop EM methods.
Xu et al. (xu1994alternative, ) demonstrated their approach on applications such as piecewise polynomial regression and classifier combination, reporting faster convergence and competitive or improved accuracy relative to earlier MoE methods. Their work provides an early theoretical foundation for quadratic gating, directly connecting to modern studies exploring quadratic MoE structures in contemporary deep learning frameworks.
The gating function $\mathcal{G}({\bm{x}})_{i}$ essentially models the posterior probability $\mathbb{P}(\zeta=i\mid{\bm{x}})$, indicating the likelihood that ${\bm{x}}$ is assigned to the partition associated with the $i$-th expert. Here, $\zeta\in\{1,\ldots,N\}$ is a latent gating variable that selects a particular expert. More precisely, the gating function defined in Equation (26) interprets this posterior probability when $\zeta$ follows a categorical distribution with parameters $(\pi_{1},\ldots,\pi_{N})^{\top}\in\Delta^{N-1}$, and conditioned on the event that $\zeta$ selects the $i$-th expert, the distribution of ${\bm{x}}$ is modeled by a specific parametric distribution $p(~{}\cdot\mid{\bm{\theta}}_{i})$. Motivated by this interpretation of gating function, Liao et al. (liao2007quadratically, ) introduced Quadratically Gated Mixture of Experts (QGME), a statistical model for multi-class, nonlinear classification when input features are partially missing.
Theoretical results for MoE models. Previous works have theoretically investigated MoE models in two main aspects, including the statistical convergence behavior of MoE, and the applications of MoE in deep learning. Regarding the statistical convergence behavior of MoE, Mendes et al. mendes2011convergence took into account a mixture of polynomial regression experts and studied the maximum likelihood estimator (MLE) of that model. In particular, they established the convergence rate of the Kullback-Leibler (KL) divergence between the estimated density and the true density as the sample size increased, and then offered useful insight into the trade-off between the model complexity and the number of experts based on these results. Next, Ho et al. ho2022gaussian and Nguyen et al. nguyen2023demystifying continued to investigate the MLE convergence behavior but for Gaussian MoE models with covariate-free and covariate-dependent gating functions, respectively, where the response variable conditioning on the covaritate variable followed from a mixture of Gaussian distributions. Their results indicated that the convergence rates of parameter estimations might be substantially slowed down due to parameter interactions expressed via some partial differential equations (PDEs). Unlike these works, Nguyen et al. nguyen2024leastsquare then considered a regression framework where the regression function admitted the form of a softmax gating MoE, and then analyzed the convergence of least squares estimators of parameters and experts. They observed that feed-forward expert networks have much faster estimation rates than polynomial experts.
Regarding the applications of MoE in deep learning, Chen et al. chen2022theory demonstrated that a single expert was incapable of learning a classification problem with intrinsic cluster structures. On the other hand, this problem would be addressed by using a sparsely gated MoE with each expert formulated as a two-layer non-linear convolutional neural networks. Next, MoE was shown to effectively solve the catastrophic forgetting issue in continual learning through overparameterized linear regression tasks li2025cl ; le2024mixture . More specifically, they illustrated that experts can be trained to specialize in various tasks, while the router can choose appropriate experts for each task. Lastly, MoE and its hierarchical variant with Laplace gating function were also theoretically verified to perform well in multimodal learning where the data included different data modalities, namely text, images, time series, and tabular data in han2024fusemoe and nguyen2024hmoe , respectively.
5 Experiments
In this section, we conduct several experiments to verify: (i) the theoretical results presented in Section 3, (ii) the empirical benefits of quadratic gating over standard linear gating in language modeling, and (iii) the favorable performance of the proposed active-attention mechanism (25) over standard attention mechanism. Furthermore, we also perform an ablation study on the number of heads when using the proposed active-attenion mechanism, which is deferred to Appendix E due to the space limit. Meanwhile, all the experimental details are provided in Appendix C.
5.1 Numerical Experiments
We generate synthetic data based on the model described in equation (9).
Results. As shown in Figure 1, the empirical convergence rates of the Voronoi losses follow the trends predicted by our theory in Section 3. For the quadratic monomial gate (Figure 1a), the loss for ReLU experts decreases at a rate close to $\mathcal{O}(n^{-0.52})$, while the loss for linear experts shows a much slower rate of around $\mathcal{O}(n^{-0.09})$. For the quadratic polynomial gate (Figure 1b), we again observe a rate of about $\mathcal{O}(n^{-0.49})$ for ReLU experts and a slower rate of $\mathcal{O}(n^{-0.06})$ for linear experts. These results are consistent with our theoretical analysis.
@@ 5.1 Numerical Experiments
5.2 Quadratic Gating versus Linear Gating
To assess the impact of gating function expressiveness on downstream performance, we conduct a set of experiments comparing quadratic and linear gating in GPT2-based MoE models.
Experimental setup. We evaluate the impact of gating function design in MoE models using the GPT2 (124M) architecture (radford2019language, ). All MoE models use 8 experts with Top2 routing. We compare three configurations: a standard dense GPT2 model, a GPT2-MoE model with a linear gating network, and a GPT2-MoE model with a quadratic gating network. To ensure fairness, we control for the total number of activated parameters across models by adjusting the hidden sizes of the experts accordingly. All models are trained on a 10 billion token subset of the FineWeb-Edu dataset (penedo2024fineweb, ).
Results. Table 3 summarizes the performance of each model on validation loss and the HellaSwag (zellers2019hellaswag, ) benchmark. The MoE model with a quadratic gating function consistently achieves the best results across both metrics. It shows improved generalization compared to the dense GPT2 baseline and demonstrates more effective expert selection than the linear gating variant. While the linear MoE slightly improves over the dense model in terms of training loss, it underperforms on HellaSwag, suggesting limited transferability. These results highlight the advantage of using a more expressive gating function in downstream tasks.
  Model Val. Loss HellaSwag (%)
Dense 3.0211 30.74%
MoE (linear) 2.9928 29.95%
MoE (quadratic) 2.9712 32.11%
5.3 Performance of Active-Attention Mechanism
We empirically demonstrate the effects of employing nonlinear activation functions in the proposed active-attention mechanism (25). Our experiments span large-scale image classification tasks on CIFAR-10 (krizhevsky2009learning, ) and ImageNet (russakovsky2015imagenet, ), language modeling on WikiText-103 (merity2016pointer, ), and multivariate time series forecasting across 8 benchmarks. We evaluate the impact of 5 commonly used activation functions, including ReLU (agarap2018deep, ), GELU (hendrycks2016gaussian, ), SiLU (elfwing2018sigmoid, ), Sigmoid, and Tanh (nwankpa2018activation, ). Figures 2 (a) and (b) present the results of image classification using different activation functions in self-attention, with ViT (dosovitskiy2020image, ) and CaiT (touvron2021going, ) as the base models. For CIFAR-10, we employed the ViT-Tiny and CaiT-Tiny models, while for ImageNet, we utilized the ViT-Base and CaiT-Medium models. Figure 2 (c) displays the results of the language modeling task on WikiText-103, using the standard multi-head self-attention transformer (vaswani2017attention, ). Additionally, we tested various activation functions on the Performer model (choromanski2020rethinking, ) as another backbone. Our findings show that the GELU and ReLU activation functions greatly improve performance compared to linear activation functions. This aligns with prior research, suggesting that these two activation functions are preferred in large-scale deep networks due to their ability to support more efficient and stable training.
@@ 5.3 Performance of Active-Attention Mechanism
Table 4 further evaluates the impact of different activation functions on transformer-based time-series forecasting models across 8 forecasting tasks. In this experiment, we employ the state-of-the-art PatchTST model (Yuqietal-2023-PatchTST, ) and the standard self-attention transformer as the backbone. Unlike the results observed in Figure 2, in addition to GELU, we find that Tanh and Sigmoid functions also show prominent advantages over linear activation function. We speculate that this is due to the smoothing gradients provided by Tanh and Sigmoid, which may help in capturing subtle patterns in time-series data. Additionally, since these tasks tend to be smaller and more prone to overfitting, the saturation effects of Tanh and Sigmoid could serve as a regularization mechanism by limiting output ranges and avoiding extreme activations.
@@ 6 Discussion
In this paper, we first establish a link between the MoE framework and the attention mechanism by showing that each row of a self-attention matrix can be interpreted as a quadratic gating MoE. Inspired by this relation, we carry out the convergence analysis of parameter estimation and expert estimation under MoE models with the quadratic polynomial gate and the quadratic monomial gate, respectively. Our theories indicate that experts formulated as neural networks with non-linear activation functions have significantly faster estimation rates than linear experts. Based on this theoretical insight on MoE and the above connection between MoE and self-attention, we propose a new attention mechanism termed active-attention, where we apply a non-linear activation function to the value matrix. Through extensive real-world experiments, we empirically demonstrate that the proposed active-attention enjoys preferable performance to the standard self-attention in various tasks, including image classification, language modeling, and multivariate time series forecasting.
Limitations. There are some limitations of our work. Firstly, in the convergence analysis of quadratic gating MoE, we take into account a single MoE layer rather than multiple layers as in practice dai2024deepseekmoe ; jiang2024mixtral for the sake of theory. Secondly, as we keep ground-truth parameters in that analysis independent of the sample size, the parameter and expert estimation rates are pointwise but not uniform. However, our results can be naturally extended using the techniques of deriving minimax optimal rates for parameter estimation in MoE from yan2025contaminated . As these two problems are not the main focus of our work, we leave them for future development.
Future directions. A potential direction of extending our theories is to investigate quadratic gating MoE in the context of a classification problem as considered in chen2022theory ; nguyen2024general . Meanwhile, it is of great importance to demystify the expressive power of quadratic gating MoE wang2025expressivepower , which has remained elusive in the literature.
@@
Supplement to “Quadratic Gating Mixture of Experts: Statistical Insights into Self-Attention”
In this supplementary material, we first provide rigorous proofs for the theoretical results of Section 3 in Appendix A. Next, we study the identifiability of the quadratic gating MoE in Appendix B. In Appendix C, we specify the details of experiments exhibited in Section 5 before discussing the computational overheads of the quadratic gating functions in Appendix D. Lastly, we carry out an empirical ablation study on the number of attention heads when using the proposed active-attention mechanism in Appendix E.
Appendix A Proof of Theoretical Results